The mass ofa nonvolatile solute of malar mass 40g/mol that should be dissolved in114g of octane to lower its vapour press.by 20% is.........
 

Grade:12th pass

1 Answers

Vikas TU
7 years ago
For immiscible mixtures, the relation in pre partial pressures and mole fractioon is:
Pa’/Pb’  = na/nb
0.2P/P = na/(114/114)
na = 0.2 moles of solute.
Mass of the solute = 0.2 * 40 => 8 grams.